OpenStudy (anonymous):

Elimination would be the most efficient way to solve which of the following systems? a. 4x+3y=6 3y=5 b. 2x+3y=12 5x-5y=21 c. y=2x x+3y=5 d. x+y=7 x-y=12

OpenStudy (mathmale):

Think: What does "elimination by addition/subtraction" mean?

OpenStudy (anonymous):

i really dont know what that means but i think its b.

OpenStudy (mathmale):

It'd be far more efficient (and produce higher grades) to learn this and other methods instead of guessing which possible answer is the correct one. suppose I were to write out a system of linear equations as x + y = 7 x - y = 12 Simply add together all the x terms, all the y terms, and all the constant terms. What do you get?

OpenStudy (anonymous):

can you explain it because im a hard learner or just give me the answer

OpenStudy (mathmale):

This material is well within your ability to learn. x+x = ? y - y =? 7+12 = ?

OpenStudy (anonymous):

2x no y 19

OpenStudy (mathmale):

Right. Therefore, x + y = 7 x - y = 12 --------- 2x = 19 Solve this for x.

OpenStudy (anonymous):

thank you so it d.

OpenStudy (mathmale):

Yes.